OpenStudy (anonymous):

Find each probability if 13 cards are drawn from a standard deck of cards and no replacement occurs.

OpenStudy (anonymous):

1. P(all hearts) 2. P(all red cards) 3. P(all one suit) 4. P(no kings)

OpenStudy (anonymous):

For 1: There are 13 cards in the deck. So how many different ways can we choose 13 cards from the set of 13 hearts? 1. And how many different ways can we choose 13 cards from the deck of 52? \(52 \choose 3\). Therefore the probability of choosing 13 hearts at random from the deck is: \[\frac{1}{52 \choose 3} = \frac{1}{\frac{52!}{3!49!}}= \frac{1}{22100}\]

OpenStudy (anonymous):

That's not right.

OpenStudy (anonymous):

Should be 52 choose 13. I dunno why I have a 3 there.

OpenStudy (anonymous):

\[\frac{1}{52 \choose 13} = \frac{1}{\frac{52!}{13!39!}}= \frac{1}{635013559600}\]

OpenStudy (anonymous):

13 cards are drawn from = (13/52)*(12/51)*(11/50) ..up to 13 ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

1 and 2 the same ?

OpenStudy (anonymous):

No, because there are twice as many red cards as there are hearts.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

(26/52)*(25/51)....up to 13 ?

OpenStudy (anonymous):

13 times. Yes

OpenStudy (anonymous):

3. P(all one suit) (52/52)*(51/51).....up to 13 =1 ?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

For P(all one suit) you have to realize it doesn't matter what you draw the first time, but each time after that you have to draw the same suit you picked first. So lets say you draw a heart first. What is the probability you'll draw a heart second?

OpenStudy (anonymous):

12/51

OpenStudy (anonymous):

And then drawing a heart the 3rd time?

OpenStudy (anonymous):

11/50

OpenStudy (anonymous):

And so on 12 times.

OpenStudy (anonymous):

than start (13/52)*(12/51)*(11/50).....up 13 the same number 1 ?

OpenStudy (anonymous):

Which one is this for?

OpenStudy (anonymous):

3. P(all one suit)

OpenStudy (anonymous):

For P(all one suit) you don't care about the first card. so it's just (12/51)*(11/50)*...*(1/40)

OpenStudy (anonymous):

4. P(no kings) 4 suite , each 13 cards , no kings 13x3= 39 stars : ( 39/52)*(38/51)......13 time

OpenStudy (anonymous):

I disagree with the 39. From each suit there is 1 king. So there are 12 cards from each suit that are not kings. Therefore there are 12*4 = 48 cards that are not kings to choose from initially: (48/52)*(47/51)*...*(36/40)

OpenStudy (anonymous):

OK,TY